A ball is thrown vertically upward from the ground with a velocity of 30m/s.

a. How long will it rise to its highest point?

b. How high does the ball rise?

c. How long after it was thrown will the ball return to the ground?


1
Expert's answer
2021-02-25T10:53:08-0500

a)


"t=\frac{v}{g}=\frac{30}{9.8}=3.06\ s"

b)


"h=\frac{v^2}{2g}=\frac{30^2}{2\cdot9.8}=45.9\ m"

c)


"T=2t=2\frac{v}{g}=2\frac{30}{9.8}=6.12\ s"